a sum of money was shared among 3 friends :albert , michael and moses in the ratio 3:5:6 . of michael received 196 more than albert find the sum of money shared ?

Answer by ikleyn(52778)

.
Thus the friends have the amounts of 3x (Albert), 5x (Michael) and 6x (Moses).


In addition, you are given that  5x - 3x = 196,


which implies  2x = 196,  x = 196/2 = 98.


From this point, please complete the solution on your own.